# Invariant nature of Raychaudhuri equation in the presence of Inflaton fields

### v1

@@ Abstract
We show that the Raychaudhuri equation remains invariant for certain solutions of scalar fields $\phi$ whose Lagrangian is non-canonical and of the form $\mathcal{L}(X,\phi)=-V(\phi)F(X)$, with $X=\frac{1}{2}g_{\mu\nu}\nabla^{\mu}\phi\nabla^{\nu}\phi$, $V(\phi)$ the potential. Solutions exist for both homogeneous and inhomogeneous fields and are reminiscent of inflaton scenarios.
@@
04.20.-q, 04.20.Cv, 98.80.-k
Introduction: The Raychaudhuri equation [1, 2, 3, 4, 5, 6], is an essential component of contemporary cosmology. The equation describes the investigation and evaluation of the time-dependent characteristics of the congruence of geodesics, as witnessed by another nearby congruence of geodesics throughout its development. The evolution primarily describes the progression of congruence inside a specific spacetime framework, and this progression is produced by a curve that follows the tangent vectors (which can be either timelike or null). The equation describes the rate at which the expansion of the congruence (a measurement that is not reliant on coordinates) is related to the gravitational influence. The equation was derived by Raychaudhuri solely based on geometrical concepts, without any reliance on other gravitational theories. The Raychaudhuri equation is commonly utilized to illustrate the attractive nature of gravity, which exerts a force between two material things. Subsequently, it emerged as a crucial lemma for investigating the renowned Hawking-Penrose singularity theorem[7, 8, 9, 10]. On a basic level, the Raychaudhuri equation takes into account the non-isotropy (regarding rotation) and non-homogeneity (regarding shear) of the cosmos [6]. The equation gives the expansion rates of two null or timelike geodesics, which can converge or diverge depending on the energy conditions provided by the Ricci tensor. Converging geodesics can be used to explain the focusing theorem [11, 12], while diverging geodesics can help to understand the non-focusing theorem due to the expansion of the universe [12, 13].
To explore many world possibilities, Das et al. [12] investigated the Raychaudhuri equation in K-essence geometry, a non-canonical theory. The authors showed that the modified Raychaudhuri equation allows for both the conditional singular and non-singular structure of the universe for cases where the spacetime is of the flat K-essence emergent Friedmann-Lemaitre-Robertson-Walker (FLRW) type. They also dealt with the development of conditional caustic universes and the Focusing theorem. An example of a non-canonical theory that investigates several avenues in cosmological research is the K-essence theory [12, 20, 14, 15, 17, 18, 19, 21, 23, 24, 25, 26, 16, 22]. There are different ways to write the Lagrangian of the K-essence field. One way is $\mathcal{L}(X,\phi)=-V(\phi)F(X)$ [20, 14, 15, 17, 18, 19], another is $\mathcal{L}(X,\phi)=F(X)-V(\phi)$ [24, 25], and a third is $\mathcal{L}(X,\phi)\equiv\mathcal{L}(X)=F(X)$ [23, 26]. Here, $F(X)\equiv\mathcal{L}(X)(\neq X)$ is the non-canonical kinetic part with $X=\frac{1}{2}g_{\mu\nu}\nabla^{\mu}\phi\nabla^{\nu}\phi$, $V(\phi)$ is the canonical potential part. This theory emphasizes that the kinetic element dominates the potential part and shows a dynamic behavior that assures late-time acceleration without fine-tuning. Another noteworthy part of the K-essence idea is its capacity to create dark energy with a sound speed ($c_{s}$) continuously lower than light. This property may lessen cosmic microwave background disruptions at large angular scales. In addition, according to Scherrer [23], this theory may be applied to both dark energy and dark matter to form a unified theory. The Dirac-Born-Infeld (DBI) type non-canonical action [27, 28] was utilized by Manna et al. [29, 30, 31, 32, 33, 34, 35, 36], and they have showed that the K-essence theory can be applied to the models of dark energy [29, 30, 34] as well as in gravitational perspectives [31, 32, 35, 36], without considering the dark sectors of the universe. The results from the Planck collaborations [37, 38, 39] have investigated the empirical evidence that supports the idea of K-essence with a DBI-type non-canonical Lagrangian among other things. The dynamical solutions of the K-essence equation of motion distinguish K-essence theories with non-canonical kinetic terms from relativistic field theories with canonical terms. These demonstrate spontaneous Lorentz invariance breakdown and metric changes for perturbations. The K-essence field Lagrangian’s theoretical formulation has non-canonical features that the metric demands. These features distribute perturbations across the evolving analogue spacetime, or curved spacetime. This can imitate the background metric, creating late-time acceleration at the right time frame in the universe’s evolution, which answers the ’cosmic coincidence’ problem.
@@
This letter aims to demonstrate the invariance of the Raychaudhuri equation in the presence of both homogeneous and inhomogeneous scalar fields within the framework of K-essence geometry. The aforementioned fields exhibit conditional correspondences with inflaton fields, which persist over the whole of the universe’s early to late phases.
K-essence: The action of the K-essence geometry can be written as [14, 15, 16, 20, 23]
@@
The energy-momentum tensor is defined as [16, 20]:
@@
We may formulate a scalar field equation of motion (EOM) in accordance with [16, 20, 29, 30]
@@
$\tilde{G}^{\mu\nu}\equiv\frac{c_{s}}{\mathcal{L}_{X}^{2}}[\mathcal{L}_{X}g^{\mu\nu}+\mathcal{L}_{XX}\nabla^{\mu}\phi\nabla^{\nu}\phi]$ (4)
with $(1+\frac{2X\mathcal{L}_{XX}}{L_{X}})>0$ and $c_{s}^{2}(X,\phi)\equiv{(1+2X\frac{\mathcal{L}_{XX}}{\mathcal{L}_{X}})^{-1}}$.
After a conformal transformation [29, 30, 31, 32] $\bar{G}_{\mu\nu}\equiv\frac{c_{s}}{\mathcal{L}_{X}}G_{\mu\nu}$, we can write the inverse metric of Eq. (4) as
@@
The corresponding geodesic equation can also be changed as [29, 30, 31, 32]
@@
where $\lambda$ is an affine parameter and the new connection coefficients ($\bar{\Gamma}^{\alpha}_{\mu\nu}$) can be expressed as
$\bar{\Gamma}^{\alpha}_{\mu\nu}=\Gamma^{\alpha}_{\mu\nu}-\frac{1}{2(1-2X)}\Big{[}\delta^{\alpha}_{\mu}\partial_{\nu}+\delta^{\alpha}_{\nu}\partial_{\mu}\Big{]}X,$ (7)
@@
with the inverse emergent metric $\bar{G}^{\mu\nu}$ defined through $\bar{G}_{\mu\lambda}\bar{G}^{\lambda\nu}=\delta^{\nu}_{\mu}$.
Modified Raychaudhuri equation: In K-essence geometry, the Raychaudhuri equation may be expressed using the covariant derivative equation (8) for a time-like vector field $v^{\beta}$ connected to an emergent metric tensor $\bar{G}_{\mu\nu}$ as [12]
@@
where $\bar{s}$ is an affine parameter, $\tilde{\Theta}\equiv D_{\alpha}v^{\beta}$ is the scalar expansion, and $\bar{R}_{\gamma\beta}$ is the Ricci tensor of this geometry. The new geometry is torsion-free and hence the above equation (9) is generated from the commutation relationship of the covariant derivatives (8) as:
$[D_{\alpha}\,,D_{\beta}]v^{\mu}=\bar{R}^{\mu}_{\gamma\alpha\beta}v^{\gamma},$ (10)
where $\bar{R}^{\mu}_{\gamma\alpha\beta}$ is the Riemann tensor associated with the K-essence emergent spacetime. Following [12], we rewrite the modified Raychaudhuri equation of the K-essence geometry as
@@
$\tilde{\Theta}=D_{\alpha}v^{\alpha}\equiv\frac{1}{\sqrt{-\bar{G}}}\leavevmode\nobreak\ \partial_{\alpha}(\sqrt{-\bar{G}}\leavevmode\nobreak\ v^{\alpha})=3\frac{\dot{a}}{a}-\frac{v^{\mu}\partial_{\mu}X}{1-2X}.$ (12)
where $\theta=\nabla_{\alpha}v^{\alpha}$ is the usual scalar expansion, symmetric shear $\sigma_{\alpha\beta}=\frac{1}{2}\Big{(}\nabla_{\beta}v_{\alpha}+\nabla_{\alpha}v_{\beta}\Big{)}-\frac{1}{3}\theta h_{\alpha\beta}$, antisymmetric rotation $\omega_{\alpha\beta}=\frac{1}{2}\Big{(}\nabla_{\beta}v_{\alpha}-\nabla_{\alpha}v_{\beta}\Big{)}$, $2\sigma^{2}=\sigma_{\alpha\beta}\sigma^{\alpha\beta}$, $2\omega^{2}=\omega_{\alpha\beta}\omega^{\alpha\beta}$ and $\nabla_{\beta}v_{\alpha}=\sigma_{\alpha\beta}+\omega_{\alpha\beta}+\frac{1}{3}\theta h_{\alpha\beta}$ with the three dimensional hypersurface metric $h_{\alpha\beta}=g_{\alpha\beta}-v_{\alpha}v_{\beta}$. It should also be noted that in [12], the authors employed a DBI type non-canonical Lagrangian represented as $\mathcal{L}(X,\phi)=1-V\sqrt{1-2X}$, where $V(\phi)\equiv V$ is a constant potential that is significantly smaller than the kinetic energy of the K-essence scalar field. Here, in Eq. (12), the background gravitational metric is assumed to be flat FLRW. Note that Eq. (11) reduces to the original Raychaudhuri equation in regular geometry when $\phi$ is absent. Moreover, it is evident from equations (11) and (12) that when $v^{\mu}\partial_{\mu}X=0$, the original Raychaudhuri equation and the original scalar expansion are recovered [5, 6, 11]. When $v^{\mu}\partial_{\mu}X=0$ we have intriguing physical possibilities involving specific solutions of the K-essence scalar field. We now show that $v^{\mu}\partial_{\mu}X=0$ admits non-trivial solutions for $\phi$ that ensure the recovery of the original Raychaudhuri equation.
Solutions: Considering the K-essence scalar field as inhomogeneous and isotropic, we define $\phi(x^{i})\equiv\phi(r,t)$ with a flat FLRW type background gravitational metric. The expression of $X$ is given by
@@
$v^{\mu}\partial_{\mu}X=\phi_{t}\phi_{tt}-\frac{\dot{a}}{a}\frac{1}{a^{2}}\phi_{r}^{2}+\frac{\phi_{r}\phi_{rt}}{a^{2}}-\phi_{t}\phi_{tr}+\frac{1}{a^{2}}\phi_{r}\phi_{rr}.$ (14)
where $\phi_{t}=\frac{\partial\phi}{\partial t}$, $\phi_{r}=\frac{\partial\phi}{\partial r}$, $\phi_{tr}=\frac{\partial^{2}\phi}{\partial t\partial r}=\phi_{rt}$, $\phi_{tt}=\frac{\partial^{2}\phi}{\partial t^{2}}$, $\phi_{rr}=\frac{\partial^{2}\phi}{\partial r^{2}}$ and $\dot{a}=\frac{\partial a}{\partial t}$. It is important to note that the four-vector $v^{\mu}$ can be expressed as $(1,1,0,0)$ for an inhomogeneous field. However, it is crucial to bear in mind that this selection of $v^{\mu}$ represents the four-vector at a particular location in spacetime, and it may not be consistent at other locations unless there is a specific situation where this specific choice is justified. Usually, in an inhomogeneous field, the components of the four-velocity will differ at different locations, and it is necessary to define their variations over the relevant region of spacetime. The vector $v^{\mu}=(1,1,0,0)$ might be seen as a specific instance in this situation, however, it does not serve as the overall answer for an inhomogeneous field. Furthermore, it is essential to note that our study focused solely on the invariance property of the Raycahudhuri equation when scalar fields are present, without addressing the cosmological behavior of the universe. For solving the constraint equation $v^{\mu}\partial_{\mu}X=0$ we consider the K-essence scalar field as
@@
Here, $H_{0}$ represents the present value of the Hubble parameter. This specific scale factor can be utilized to describe the acceleration of the universe in later periods, when it is primarily influenced by dark energy [40], as well as during inflation in the early stages of the universe [41, 42, 43]. Therefore, the equation (16) can be separated as
@@
where $w^{2}$ is the separation constant. It’s important to mention that the separation constant $w^{2}$ is either zero or a positive non-zero value.
Case-I $w^{2}=0$: From Eq. (18) we get the solutions as ($e^{2H_{0}t}\neq 0$):
@@
where $A,\leavevmode\nobreak\ B\leavevmode\nobreak\ C\leavevmode\nobreak\ D,\leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ E$ are the arbitrary integration constants. In the first example, $\phi(r,t)$ has four distinct kinds of solutions, and $v^{\mu}\partial_{\mu}X$ is zero for all of them. The modified Raychaudhuri equation (11) reduces to the standard Raychaudhuri equation for these four categories of scalar fields. The constant scalar field is the first of the four solutions (19), the homogenous field is the second (20), and the inhomogeneous scalar fields are the last two (21 and 22). Thus, it can be concluded that the Raychaudhuri equation (11) and the scalar expansion (12) remain invariant for both the homogeneous and inhomogeneous scalar field solutions given above. Note that, the fourth solution (22) is the general solution of Eq.(18) for $w^{2}=0$, which incorporates both the space part ($r$) and the time part ($t$). For $t=0$ one can achieve the third solution (21), for $r=0$, one would get the second solution (20) and for both $t=0\leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ r=0$ one would get the first solution (19). Typically, scalar field models are employed to elucidate the inflationary phases of the universe, which refer to the rapid expansion of space within a short period. In our fourth solution (22), it is evident that the scalar field exhibits exponential dependence on space and linear dependence on time, suggesting inflationary behavior [43].
Case-II $w^{2}\neq 0$: From the Eq. (18) the general solutions of the K-essence scalar field are:
$\phi^{\pm}(r,t)$ $=\mp\frac{1}{2H_{0}^{3/2}}\Bigg{[}2\sqrt{w^{2}+e^{2H_{0}(r+G_{1})}}$ (23)
@@
$\phi^{\pm}(r,t)$ =±$=\pm\sqrt{\frac{w^{2}}{H_{0}}}\leavevmode\nobreak\ r+G_{3}\mp\frac{1}{H_{0}}\sqrt{\frac{w^{2}}{H_{0}}}e^{-H_{0}t}+F_{2}$ (24)
@@
Matching to inflaton fields: Now let’s examine the behaviors of the solutions of the fields concerning the existing inflationary fields, using a standard technique described in Mukhanov’s well-known book [43].
Case-I: Homogeneous field: Following Mukhanov [43], the inflationary phase of the cosmos can be characterized by a scalar field known as “inflaton” in conventional theory. According to the author, the ultra-hard equation of state is characterized by the condition that the kinetic component of the scalar field ($\dot{\varphi}^{2}$) exceeds the potential ($V(\varphi)$) of the field at a homogeneous cosmology. In this scenario, the equation of state can be expressed as $p\approx+\rho$, where $\rho$ represents the standard energy density of the universe. In the ultra-hard stage, the expression for the field can be expressed by

### v2

# Form Invariance of Raychaudhuri equation in the presence of Inflaton-type fields

@@ Abstract
We show that the Raychaudhuri equation remains form invariant for certain solutions of scalar fields $\phi$ whose Lagrangian is non-canonical and of the form $\mathcal{L}(X,\phi)=-V(\phi)F(X)$, with $X=\frac{1}{2}g_{\mu\nu}\nabla^{\mu}\phi\nabla^{\nu}\phi$ and $V(\phi)$ the potential. Solutions exist for both homogeneous and inhomogeneous fields that are like inflatons. Certain recent observations indicate that the cosmos is inhomogeneous and thus our results are in sync with latest observations. So the Raychaudhuri equation can accommodate primordial inhomogeneities as well as cosmologically relevant scenarios.
@@
04.20.-q, 04.20.Cv, 04.50.Kd, 98.80.-k
I Introduction:
@@
This article aims to demonstrate the form invariance of the Raychaudhuri equation (RE)  [1, 2, 3, 4, 5, 6] in the presence of both homogeneous and inhomogeneous scalar fields whose Lagrangian is non-canonical and associated with K-essence geometry. The aforementioned fields exhibit conditional correspondences with inflaton fields, which persist over the whole of the universe’s early to late phases. So the Raychaudhuri equation may also account for primordial inhomogeneities. We focus on the Raychaudhuri equation here because this is a geometric equation that has fundamental significance without relying on any theory of gravity. Cosmological consequences have been discussed elsewhere for homogeneous fields [7].
The Raychaudhuri equation describes the time-dependent characteristics of the congruence of geodesics as witnessed by another nearby congruence of geodesics throughout its development. The evolution primarily describes the progression of congruence inside a specific spacetime framework and this progression is produced by a curve that follows the tangent vectors (which can be either timelike or null). The equation describes the rate at which the expansion of the congruence (a measurement not dependent on coordinate choices) is related to the gravitational influence. It also illustrates the attractive nature of gravity. Subsequently it emerged as a crucial lemma for investigating the renowned Hawking-Penrose singularity theorem[8, 9, 10, 11]. At a more fundamental level, the Raychaudhuri equation takes into account the anisotropy (regarding rotation) and inhomogeneity (regarding shear) of the cosmos [6]. The equation gives the expansion rates of two null or timelike geodesics, which can converge or diverge depending on the energy conditions provided by the Ricci tensor. Converging geodesics can be used to explain the focusing theorem [12, 7], while diverging geodesics can help to understand the non-focusing theorem due to the expansion of the universe [7, 13]. Burger et al. [14] have discussed the relevance and several uses of the Raychaudhuri equation beyond general relativity.
The plan of this paper is as follows: in Section 2 we briefly review the theory of K-essence which involves the non-canonical Lagrangian, Section 3 discusses the Raychaudhuri equation in the presence of K-essence, section 4 deals with the homogeneous and inhomogeneous solutions of the K-essence fields that leave the Raychaudhuri equation invariant, Section 5 relates the solutions to inflaton fields while Section 6 is the conclusion.
II K-essence
An example of a non-canonical theory that investigates several avenues in cosmological research is the K-essence theory [7, 21, 15, 16, 18, 19, 20, 22, 24, 25, 26, 27, 17, 23]. There are different ways to write the Lagrangian of the K-essence field. One way is $\mathcal{L}(X,\phi)=-V(\phi)F(X)$ [21, 15, 16, 18, 19, 20], another is $\mathcal{L}(X,\phi)=F(X)-V(\phi)$ [25, 26], and a third is $\mathcal{L}(X,\phi)\equiv\mathcal{L}(X)=F(X)$ [24, 27]. Here, $F(X)\equiv\mathcal{L}(X)(\neq X)$ is the non-canonical kinetic part with $X=\frac{1}{2}g_{\mu\nu}\nabla^{\mu}\phi\nabla^{\nu}\phi$, $V(\phi)$ is the canonical potential part. This theory emphasizes that the kinetic element dominates the potential part and shows a dynamic behavior that assures late-time acceleration without fine-tuning. Another noteworthy part of the K-essence idea is its capacity to create dark energy with a sound speed ($c_{s}$) continuously lower than light. This property may lessen cosmic microwave background disruptions at large angular scales. In addition, according to Scherrer [24], this theory may be applied to both dark energy and dark matter to form a unified theory. The Dirac-Born-Infeld (DBI) type non-canonical action [28, 29] was utilized by Manna et al. [30, 31, 32, 33, 34, 35, 36, 37], and they have shown that the K-essence theory can be applied to the models of dark energy [30, 31, 35] as well as in gravitational perspectives [32, 33, 36, 37], without considering the dark sectors of the universe. The K-essence theory has also been utilized in describing the inflationary behavior of the early universe [15, 38, 39]. The results from the Planck collaborations [40, 41, 42] have investigated the empirical evidence that supports the idea of K-essence with a DBI-type non-canonical Lagrangian, among other things. The dynamical solutions of the K-essence equation of motion distinguish K-essence theories with non-canonical kinetic terms from relativistic field theories with canonical terms. These demonstrate spontaneous Lorentz invariance breakdown and metric changes under perturbations. The K-essence field Lagrangian’s theoretical formulation has non-canonical features that the metric demands. These features distribute perturbations across the evolving analogue spacetime, or curved spacetime. This can imitate the background metric creating late-time acceleration at the right time frame in the universe’s evolution and so is relevant in the ’cosmic coincidence’ problem.
III Modified Raychaudhuri equation in the context of K-essence:
The action of the K-essence geometry can be written as [15, 16, 17, 21, 24]
@@ III Modified Raychaudhuri equation in the context of K-essence:
The energy-momentum tensor is [17, 21]:
@@
A scalar field equation of motion (EOM) is [17, 21, 30, 31]
@@
$\tilde{G}^{\mu\nu}\equiv\frac{c_{s}}{\mathcal{L}_{X}^{2}}[\mathcal{L}_{X}g^{\mu\nu}+\mathcal{L}_{XX}\nabla^{\mu}\phi\nabla^{\nu}\phi]$ with $(1+\frac{2X\mathcal{L}_{XX}}{L_{X}})>0$ and $c_{s}^{2}(X,\phi)\equiv{(1+2X\frac{\mathcal{L}_{XX}}{\mathcal{L}_{X}})^{-1}}$.
Using a conformal transformation [30, 31, 32, 33] $\bar{G}_{\mu\nu}\equiv\frac{c_{s}}{\mathcal{L}_{X}}G_{\mu\nu}$, the inverse metric of $\tilde{G}^{\mu\nu}$ is
@@
The corresponding geodesic equation is [30, 31, 32, 33]
@@
where $\lambda$ is an affine parameter and the new connection coefficients ($\bar{\Gamma}^{\alpha}_{\mu\nu}$) is
$\bar{\Gamma}^{\alpha}_{\mu\nu}=\Gamma^{\alpha}_{\mu\nu}-\frac{1}{2(1-2X)}\Big{[}\delta^{\alpha}_{\mu}\partial_{\nu}+\delta^{\alpha}_{\nu}\partial_{\mu}\Big{]}X.$ (6)
@@
The corresponding Einstein equation in this geometry is [7, 21, 43]: $\bar{\mathcal{G}}_{\mu\nu}=\bar{R}_{\mu\nu}-\frac{1}{2}\bar{G}_{\mu\nu}\bar{R}=\kappa\bar{T}_{\mu\nu}$, where $\kappa=8\pi G$ is constant and $\bar{R}_{\mu\nu}$ is the Ricci tensor, $\bar{R}\leavevmode\nobreak\ (=\bar{R}_{\mu\nu}\bar{G}^{\mu\nu})$ and $\bar{T}_{\mu\nu}$ is the energy-momentum tensor of this geometry.
In K-essence geometry, the Raychaudhuri equation for a time-like vector field $v^{\beta}$ is [7]
@@
with $\bar{s}$ is an affine parameter, $\tilde{\Theta}\equiv D_{\alpha}v^{\beta}$ the scalar expansion, and $\bar{R}_{\gamma\beta}$ the Ricci tensor of this geometry. The new geometry is torsion-free and hence equation (8) is generated from the commutation relationship of the covariant derivatives (7) as
$[D_{\alpha}\,,D_{\beta}]v^{\mu}=\bar{R}^{\mu}_{\gamma\alpha\beta}v^{\gamma}$ (9)
where $\bar{R}^{\mu}_{\gamma\alpha\beta}$ is the Riemann tensor. Following [7], we rewrite the modified Raychaudhuri equation as
@@
$\tilde{\Theta}=D_{\alpha}v^{\alpha}\equiv\frac{1}{\sqrt{-\bar{G}}}\leavevmode\nobreak\ \partial_{\alpha}(\sqrt{-\bar{G}}\leavevmode\nobreak\ v^{\alpha})=3\frac{\dot{a}}{a}-\frac{v^{\mu}\partial_{\mu}X}{1-2X}.$
where $\theta=\nabla_{\alpha}v^{\alpha}$ is the usual scalar expansion, symmetric shear $\sigma_{\alpha\beta}=\frac{1}{2}\Big{(}\nabla_{\beta}v_{\alpha}+\nabla_{\alpha}v_{\beta}\Big{)}-\frac{1}{3}\theta h_{\alpha\beta}$, antisymmetric rotation $\omega_{\alpha\beta}=\frac{1}{2}\Big{(}\nabla_{\beta}v_{\alpha}-\nabla_{\alpha}v_{\beta}\Big{)}$, $2\sigma^{2}=\sigma_{\alpha\beta}\sigma^{\alpha\beta}$, $2\omega^{2}=\omega_{\alpha\beta}\omega^{\alpha\beta}$ and $\nabla_{\beta}v_{\alpha}=\sigma_{\alpha\beta}+\omega_{\alpha\beta}+\frac{1}{3}\theta h_{\alpha\beta}$ with the three dimensional hypersurface metric $h_{\alpha\beta}=g_{\alpha\beta}-v_{\alpha}v_{\beta}$. In [7], the authors employed a DBI type non-canonical Lagrangian represented as $\mathcal{L}(X,\phi)=1-V\sqrt{1-2X}$, where $V(\phi)\equiv V$ is a constant potential that is significantly smaller than the kinetic energy of the K-essence scalar field. Here, in Eq. (11), the background gravitational metric is assumed to be flat FLRW.
Note that Eq. (10) is the Raychaudhuri equation in the new geometry when $\phi$ is present. This is evident from equations (10) and (11) that this modified RE is form invariant when $v^{\mu}\partial_{\mu}X=0$ [5, 6, 12]. The modified scalar expansion reduces to its original form. It is important to note that the Ricci tensor ($\bar{R}_{\gamma\beta}$) on the right side of the modified RE (10) differs from the original Ricci tensor ($R_{\gamma\beta}$) which defines a different geometry. This is as per the dynamical behavior of spacetime. However, we know that the Raychaudhuri equation is an identity (independent of any gravitational theories, created geometrically) and therefore form invariance of RE should be a logical consequence in any geometry defined by the Ricci tensor. We now show that $v^{\mu}\partial_{\mu}X=0$ admits non-trivial solutions for $\phi$ that ensure the recovery of the form of the original Raychaudhuri equation. Moreover, these solutions of $\phi$ are similar to inflatons, which are cosmologically significant in the early universe. Note that we have obtained cosmologically relevant solutions without taking direct recourse to the Friedmann equations, which is the usual practice.
IV Solutions:
Considering the K-essence scalar field as inhomogeneous and isotropic, we define $\phi(x^{i})\equiv\phi(r,t)$ with a flat FLRW type background gravitational metric, $X$ is
@@ IV Solutions:
$v^{\mu}\partial_{\mu}X=-\phi_{t}\phi_{tt}-\frac{\dot{a}}{a}\frac{1}{a^{2}}\phi_{r}^{2}+\frac{\phi_{r}\phi_{rt}}{a^{2}}-\phi_{t}\phi_{tr}+\frac{1}{a^{2}}\phi_{r}\phi_{rr}.$
where $\phi_{t}=\frac{\partial\phi}{\partial t}$, $\phi_{r}=\frac{\partial\phi}{\partial r}$, $\phi_{tr}=\frac{\partial^{2}\phi}{\partial t\partial r}=\phi_{rt}$, $\phi_{tt}=\frac{\partial^{2}\phi}{\partial t^{2}}$, $\phi_{rr}=\frac{\partial^{2}\phi}{\partial r^{2}}$ and $\dot{a}=\frac{\partial a}{\partial t}$. The four vectors, $v^{\mu}$ may be expressed as $(1,1,0,0)$ for an inhomogeneous field. However, this selection is for a particular location in spacetime and may not be consistent elsewhere unless this specific choice is justified. $v^{\mu}=(1,1,0,0)$ might be seen as a specific instance in this situation, however, it does not serve as the overall answer for an inhomogeneous field. Furthermore, our study focuses solely on the invariance property of the Raychaudhuri equation when scalar fields are present without addressing any specific cosmological situation. Therefore, obtaining any result that may have any cosmological significance is a bonus. To solve $v^{\mu}\partial_{\mu}X=0$ take $\phi$ as
@@
$H_{0}$ represents the positive constant value of Hubble parameter which may vary in different epochs. The specific scale factor can be utilized to describe the acceleration of the universe in later periods when it is primarily influenced by dark energy [44] as well as during inflation in the early universe [45, 46, 38]. Basically during inflation the time derivative of the scale factor ($da(t)/dt$) was positive, in this case which emerges as $\dot{a}/a=H_{0}$. Therefore, the equation (15) becomes
@@
where $w^{2}\leavevmode\nobreak\ (\geq 0)$ is the separation constant.
Case-I: $w^{2}=0$: From Eq. (17) we get the solutions as ($e^{2H_{0}t}\neq 0$):
@@
where $A,\leavevmode\nobreak\ B,\leavevmode\nobreak\ C,\leavevmode\nobreak\ D\leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ E$ are the arbitrary integration constants. In the first example, $\phi(r,t)$ has four distinct kinds of solutions, and $v^{\mu}\partial_{\mu}X$ is zero for all of them. The modified RE (10) reduces to the standard RE for these four categories of scalar fields. The constant scalar field is the first of the four solutions (18), the homogeneous field is the second (19), and the inhomogeneous scalar fields are the last two (20 and 21). Thus, it can be concluded that the modified RE (10) and the scalar expansion (11) remain invariant for both the homogeneous and inhomogeneous scalar field solutions given above. Note that the fourth solution (21) is the general solution of Eq.(17) for $w^{2}=0$, which incorporates both the space part ($r$) and the time part ($t$). For $t=0$ one can achieve the third solution (20), for $r=0$, one would get the second solution (19) and for both $t=0\leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ r=0$ one would get the first solution (18). Typically, scalar field models are employed to elucidate the inflationary phases of the universe, which refer to the rapid expansion of space within a short period. In our fourth solution (21), it is evident that the scalar field exhibits exponential dependence on space and linear dependence on time, suggesting inflationary behavior [38].
Case-II: $w^{2}\neq 0$: From the Eq. (17) the general solutions of the K-essence scalar field are:
$\phi^{\pm}(r,t)$ $=\mp\frac{1}{2H_{0}^{3/2}}\Bigg{[}2\sqrt{w^{2}+e^{2H_{0}(r+G_{1})}}$
@@
$\phi^{\pm}(r,t)=\pm\sqrt{\frac{w^{2}}{H_{0}}}\leavevmode\nobreak\ r+G_{3}\mp\frac{1}{H_{0}}\sqrt{\frac{w^{2}}{H_{0}}}e^{-H_{0}t}+F_{2}$ (23)
@@
V Inflaton fields:
We now show that our solutions (19), (21) and (22) are similar to inflaton fields as described in Mukhanov [38].
Case-I: Homogeneous field: The inflationary phase of the cosmos can be characterized by a scalar field known as “inflaton”. In the ultra-hard equation of state the kinetic component of the scalar field $(\dot{\varphi}^{2})>>V(\varphi)$, where $V(\varphi)$ is the potential term, and the equation of state is $p\approx+\rho$, where $\rho$ represents the standard energy density of the universe. In the ultra-hard stage [38]